OpenStudy (lilai3):

Hey! Who wuvzzzzzzz money??? $$$$$ meeee Deb is saving $54 per month to buy a new camera. Use the Distributive Property to mentally find how much she has saved after 7 months. Explain.

OpenStudy (hoblos):

each month she saves $54 so in 7 months she saves 7(54) = $378

OpenStudy (lilai3):

okay thank you hobolos

OpenStudy (hoblos):

glad to help :)

OpenStudy (lilai3):

wait but ur suppose to use the distributive property

OpenStudy (lilai3):

is it 12 (54+7) ????

OpenStudy (anonymous):

no

OpenStudy (anonymous):

thatll yield a whole different number

OpenStudy (lilai3):

oh kay so then wut's da correct answer?

OpenStudy (hoblos):

well i cant find another way than what i posted!! i think thats it just 7(54)

OpenStudy (anonymous):

7(54) = $378

OpenStudy (lilai3):

but it is suppose to be the distributive property way it says in problem

OpenStudy (anonymous):

if you really wanna distribute you can do this 7(a+b+c) where a b c ad up to 54 :P

OpenStudy (lilai3):

what?000000000

OpenStudy (lilai3):

oops

OpenStudy (anonymous):

7(50+2+2)=350+14+14=378

OpenStudy (anonymous):

as you can see 7*5 is easy it is 35 just add a 0 to end 350. 2*7 is 14 so it is just 350+14+14

OpenStudy (lilai3):

uhm okay but i don't really gert it

OpenStudy (lilai3):

get

OpenStudy (anonymous):

lol :) well i think your problem is as simple as it gets no need for distributive property :)

OpenStudy (anonymous):

7(54) = $378 in fact is right you distributed the 7 to the 54 :)

OpenStudy (lilai3):

oh kay so do i just stick with the first one that you said that equals to 3 hundred sumthing dollars right?

OpenStudy (lilai3):

ohokay thanks so

OpenStudy (anonymous):

stick to first one :)